# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ A.3.3 Proof of Lemma 13
$=\;\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})+\theta\big{)}\big{|}$
$\leq\;\Big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big{)}\Big{|}$
$\;\quad+\Big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})+\theta\big{)}\Big{|}$
@@
where the first equality holds since, for any $\gamma\geq 0$, it can be checked that the inequality $\Delta_{\theta}=\max(0,D(\mathbb{X}_{n})-\theta)>\gamma$ holds if and only if $D(\mathbb{X}_{n})<\gamma+\theta$. We will now bound the terms $T_{1}$ and $T_{2}$, respectively. Denote for brevity $\gamma_{\mathbf{z}}\coloneqq q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta$ and $\gamma_{\ast}\coloneqq q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})+\theta$. Fix any $\delta>0$. Applying Lemma 12 with $\gamma=\gamma_{\mathbf{z}}$ yields
@@
where $t_{m,n,\delta}\coloneqq\big{(}4\sqrt{2}\tau_{\infty}\epsilon_{m,n}/\sqrt{\delta n}+2\tau_{\infty}\epsilon_{m,n}^{2}\big{)}^{1/2}$ is defined in Lemma 12, and (39) follows from a triangle inequality. Moreover, by Lemma 11, there exists an event, say $\mathcal{A}$, with probability at least $1-\delta$ such that, on $\mathcal{A}$,
@@
$\gamma_{\ast}-\rho_{m,n,\delta}\;\leq\;\gamma_{\ast}-\big{|}\gamma_{\mathbf{z}}-\gamma_{\ast}\big{|}\;\leq\;\min(\gamma_{\mathbf{z}},\;\gamma_{\ast})\;\leq\;\max(\gamma_{\mathbf{z}},\;\gamma_{\ast})\;$ $\leq\;\gamma_{\ast}+\big{|}\gamma_{\mathbf{z}}-\gamma_{\ast}\big{|}\;\leq\;\gamma_{\ast}+\rho_{m,n,\delta}\;.$
@@
$t_{m,n,\delta}\;=\;\left(4\sqrt{2}\tau_{\infty}\delta^{-\frac{1}{2}}\epsilon_{m,n}n^{-\frac{1}{2}}+2\tau_{\infty}\epsilon_{m,n}^{2}\right)^{\frac{1}{2}}\;$ $\leq\;\left(4\sqrt{2}\tau_{\infty}\delta\right)^{\frac{1}{2}}\epsilon_{m,n}^{\frac{1}{2}}n^{-\frac{1}{4}}+\sqrt{2}\tau_{\infty}^{\frac{1}{2}}\epsilon_{m,n}\;\in\;o\left(n^{-\frac{1}{2}}\right)\;,$
@@
$\rho_{m,n,\delta}\;=\;\frac{\tau_{\infty}^{\frac{1}{2}}\epsilon_{m,n}^{\frac{1}{4}}}{n^{\frac{1}{2}}}\sqrt{C_{1}+C_{2}\log\big{(}8n^{2(1-2s)}\big{)}}\;$ $=\;\frac{\tau_{\infty}^{\frac{1}{2}}\epsilon_{m,n}^{\frac{1}{4}}}{n^{\frac{1}{2}}}\sqrt{C_{1}^{\prime}+C_{2}^{\prime}\log(n)}\;\in\;o\left(n^{-\frac{1}{2}}\right)\;,$ (43)
@@
$\max_{m^{\prime}\leq f_{n}}\sup_{\mathbf{z}_{1},\ldots,\mathbf{z}_{m^{\prime}}\in\mathbb{R}^{d}}\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}$
@@
To complete our proof, it remains to show that the first term can be bounded by $\delta$, from which the claimed result in Lemma 13 would follow by redefining $\delta$. To proceed, we write
@@
$=\;\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})\;\leq\;q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})+\eta\left(f_{n}\right)\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})\;\leq\;q_{\infty,1-\alpha}(\mathbb{X}_{n})-\eta\left(f_{n}\right)\big{)}$
$=\;\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\sqrt{n}D(\mathbb{X}_{n}^{\ast})\;\leq\;\sqrt{n}\zeta_{n,\theta}+\sqrt{n}\eta\left(f_{n}\right)\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\sqrt{n}D(\mathbb{X}_{n}^{\ast})\;\leq\;\sqrt{n}\zeta_{n,\theta}-\sqrt{n}\eta\left(f_{n}\right)\big{)}$ (44)
@@
Under the assumed conditions on $k$ and assuming $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$, Gorham and Mackey, (2017, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$, so the V-statistic $D^{2}(\mathbb{X}_{n}^{\ast})$ is degenerate (see, e.g., Serfling,, 2009, Chapter 6), and classic results on the asymptotics of degenerate V-statistics (Serfling,, 2009, Theorem 6.4.1 A) shows that $nD^{2}(\mathbb{X}_{n}^{\ast})$ converges weakly to a non-negative distribution. Since the square-root function $x\mapsto\sqrt{x}$ is continuous on $[0,\infty)$ and continuous functions preserve weak limits by the Continuous Mapping Theorem (Van der Vaart,, 2000, Theorem 2.3), the squared-root statistic, $\sqrt{n}D(\mathbb{X}_{n}^{\ast})$, also converges weakly to a non-negative distribution. In particular, the quantile $q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})$, and thus $\sqrt{n}\zeta_{n,\theta}$, converge to a positive number as $n\to\infty$. Also since $\eta\left(f_{n}\right)=o(n^{-1/2})$ as argued below (43), we have $\sqrt{n}\eta\left(f_{n}\right)\to 0$. In summary, we have shown that
@@
A.3.4 Proofs for Proposition 14 and Theorem 3
Proof of Proposition 14  For each $n$ and $R\in\mathcal{P}(\mathbb{R}^{d})$, define $Q=(1-\epsilon_{n})P+\epsilon_{n}R$. Let $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ be random samples drawn from $Q$ and $P$, respectively. As argued in Section A.1, each random variable in $\mathbb{X}_{n}$ can be written as $\mathbf{X}_{i}\stackrel{{\scriptstyle d}}{{=}}(1-\xi_{i})\mathbf{X}_{i}^{\ast}+\xi_{i}\mathbf{Z}_{i}$, where $\mathbf{Z}_{i}\sim R$ and $\xi_{i}\sim\mathrm{Bernoulli}(n,\epsilon_{n})$ are independent, and =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution. Define the random variable $M^{\prime}=\sum_{i=1}^{n}\xi_{i}$, then $M^{\prime}\sim\mathrm{Binomial}(n,\epsilon_{n})$.
@@ A.3.4 Proofs for Proposition 14 and Theorem 3
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\mathcal{A}(\mathbb{X}_{n}))\;$ $\leq\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\mathcal{A}(\mathbb{X}_{n})\cap\mathcal{B})+\Pr(\mathcal{B}^{\mathsf{c}})$
@@
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\mathcal{A}(\mathbb{X}_{n}))\;$ $\geq\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\mathcal{A}(\mathbb{X}_{n})\cap\mathcal{B})$
@@
It remains to bound the two terms on the RHS of the above inequality. The second term can be bounded by noting that, since $\epsilon_{n}\leq f_{n}/n$ and $f_{n}=o(n^{1/2})$, it is clear that $t_{n}$, defined in (45), converges to 0 as $n\to\infty$, so $t_{n}\leq\delta/2$ for sufficiently large $n$. To bound the first term, since $\epsilon_{n}n+f_{n}\leq 2f_{n}=o(n^{1/2})$ and the assumptions in Lemma 13 hold, we can apply Lemma 13 to conclude that there exists $n_{0}$ such that for any $n\geq n_{0}$, we have
@@
Proof of Theorem 3  This result immediately follows from Proposition 14 by setting $\theta=0$, in which case

### v4

@@ A.3.3 Proof of Lemma˜13
$=\;\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big{)}$
$\;\qquad-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})+\theta\big{)}\big{|}$
$\leq\;\Big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big{)}$
$\;\qquad-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big{)}\Big{|}$
$\;\quad+\Big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big{)}$
$\;\qquad\;\;\;-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})+\theta\big{)}\Big{|}$
@@
where the first equality holds since, for any $\gamma\geq 0$, it can be checked that the inequality $\Delta_{\theta}=\max(0,D(\mathbb{X}_{n})-\theta)>\gamma$ holds if and only if $D(\mathbb{X}_{n})<\gamma+\theta$. We will now bound the terms $T_{1}$ and $T_{2}$, respectively. Denote for brevity $\gamma_{\mathbf{z}}\coloneqq q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta$ and $\gamma_{\ast}\coloneqq q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})+\theta$. Fix any $\delta>0$. Applying Lemma˜12 with $\gamma=\gamma_{\mathbf{z}}$ yields
@@
where $t_{m,n,\delta}\coloneqq\big{(}4\sqrt{2}\tau_{\infty}\epsilon_{m,n}/\sqrt{\delta n}+2\tau_{\infty}\epsilon_{m,n}^{2}\big{)}^{1/2}$ is defined in Lemma˜12, and (39) follows from a triangle inequality. Moreover, by Lemma˜11, there exists an event, say $\mathcal{A}$, with probability at least $1-\delta$ such that, on $\mathcal{A}$,
@@
$\gamma_{\ast}-\rho_{m,n,\delta}\;\leq\;\gamma_{\ast}-\big{|}\gamma_{\mathbf{z}}-\gamma_{\ast}\big{|}\;\leq\;\min(\gamma_{\mathbf{z}},\;\gamma_{\ast})\;\leq\;\max(\gamma_{\mathbf{z}},\;\gamma_{\ast})\;$ $\leq\;\gamma_{\ast}+\big{|}\gamma_{\mathbf{z}}-\gamma_{\ast}\big{|}$
$\leq\;\gamma_{\ast}+\rho_{m,n,\delta}\;.$
@@
$t_{m,n,\delta}\;=\;\left(4\sqrt{2}\tau_{\infty}\delta^{-\frac{1}{2}}\epsilon_{m,n}n^{-\frac{1}{2}}+2\tau_{\infty}\epsilon_{m,n}^{2}\right)^{\frac{1}{2}}\;$ $\leq\;\left(4\sqrt{2}\tau_{\infty}\delta\right)^{\frac{1}{2}}\epsilon_{m,n}^{\frac{1}{2}}n^{-\frac{1}{4}}+\sqrt{2}\tau_{\infty}^{\frac{1}{2}}\epsilon_{m,n}$
$\in\;o\left(n^{-\frac{1}{2}}\right)\;,$
@@
$\rho_{m,n,\delta}\;=\;\frac{\tau_{\infty}^{\frac{1}{2}}\epsilon_{m,n}^{\frac{1}{4}}}{n^{\frac{1}{2}}}\sqrt{C_{1}+C_{2}\log\big{(}8n^{2(1-2s)}\big{)}}\;$ $=\;\frac{\tau_{\infty}^{\frac{1}{2}}\epsilon_{m,n}^{\frac{1}{4}}}{n^{\frac{1}{2}}}\sqrt{C_{1}^{\prime}+C_{2}^{\prime}\log(n)}$
$\in\;o\left(n^{-\frac{1}{2}}\right)\;,$ (43)
@@
$\max_{m^{\prime}\leq f_{n}}\sup_{\mathbf{z}_{1},\ldots,\mathbf{z}_{m^{\prime}}\in\mathbb{R}^{d}}\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\big{)}$
$\;\qquad\qquad\qquad\qquad\;-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}$
@@
To complete our proof, it remains to show that the first term can be bounded by $\delta$, from which the claimed result in Lemma˜13 would follow by redefining $\delta$. To proceed, we write
@@
$=\;\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})\;\leq\;q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})+\eta\left(f_{n}\right)\big{)}$
$\;\quad-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})\;\leq\;q_{\infty,1-\alpha}(\mathbb{X}_{n})-\eta\left(f_{n}\right)\big{)}$
$=\;\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\sqrt{n}D(\mathbb{X}_{n}^{\ast})\;\leq\;\sqrt{n}\zeta_{n,\theta}+\sqrt{n}\eta\left(f_{n}\right)\big{)}$ (44a)
$\;\quad-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\sqrt{n}D(\mathbb{X}_{n}^{\ast})\;\leq\;\sqrt{n}\zeta_{n,\theta}-\sqrt{n}\eta\left(f_{n}\right)\big{)}$ (44b)
@@
Under the assumed conditions on $k$ and assuming $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$, Gorham and Mackey, (2017, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$, so the V-statistic $D^{2}(\mathbb{X}_{n}^{\ast})$ is degenerate (see, e.g., Serfling,, 2009, Chapter 6), and classic results on the asymptotics of degenerate V-statistics (Serfling,, 2009, Theorem 6.4.1 A) shows that $nD^{2}(\mathbb{X}_{n}^{\ast})$ converges weakly to a non-negative distribution. Since the square-root function $x\mapsto\sqrt{x}$ is continuous on $0,\infty)$ and continuous functions preserve weak limits by the Continuous Mapping Theorem (Van der Vaart,, 2000, Theorem 2.3), the squared-root statistic, $\sqrt{n}D(\mathbb{X}_{n}^{\ast})$, also converges weakly to a non-negative distribution. In particular, the scaled quantile $\sqrt{n}q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})$, and thus also $\sqrt{n}\zeta_{n,\theta}$, converge to a positive number as $n\to\infty$. Also since $\eta\left(f_{n}\right)=o(n^{-1/2})$ as argued below (43), we have $\sqrt{n}\eta\left(f_{n}\right)\to 0$. In summary, we have shown that
@@
A.3.4 Proofs for Proposition˜14 and Theorem˜3
Proof of Proposition˜14  For each $n$ and $R\in\mathcal{P}(\mathbb{R}^{d})$, define $Q=(1-\epsilon_{n})P+\epsilon_{n}R$. Let $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ be random samples drawn from $Q$ and $P$, respectively. As argued in Section˜A.1, each random variable in $\mathbb{X}_{n}$ can be written as $\mathbf{X}_{i}\stackrel{{\scriptstyle d}}{{=}}(1-\xi_{i})\mathbf{X}_{i}^{\ast}+\xi_{i}\mathbf{Z}_{i}$, where $\mathbf{Z}_{i}\sim R$ and $\xi_{i}\sim\mathrm{Bernoulli}(n,\epsilon_{n})$ are independent, and $\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution. Define the random variable $M^{\prime}=\sum_{i=1}^{n}\xi_{i}$, then $M^{\prime}\sim\mathrm{Binomial}(n,\epsilon_{n})$.
@@ A.3.4 Proofs for Proposition˜14 and Theorem˜3
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\mathcal{A}(\mathbb{X}_{n}))$
$\leq\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\mathcal{A}(\mathbb{X}_{n})\cap\mathcal{B})+\Pr(\mathcal{B}^{\mathsf{c}})$
@@
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\mathcal{A}(\mathbb{X}_{n}))$
$\geq\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\mathcal{A}(\mathbb{X}_{n})\cap\mathcal{B})$
@@
It remains to bound the two terms on the RHS of the above inequality. The second term can be bounded by noting that, since $\epsilon_{n}\leq f_{n}/n$ and $f_{n}=o(n^{1/2})$, it is clear that $t_{n}$, defined in (45), converges to 0 as $n\to\infty$, so $t_{n}\leq\delta/2$ for sufficiently large $n$. To bound the first term, since $\epsilon_{n}n+f_{n}\leq 2f_{n}=o(n^{1/2})$ and the assumptions in Lemma˜13 hold, we can apply Lemma˜13 to conclude that there exists $n_{0}$ such that for any $n\geq n_{0}$, we have
@@
Proof of Theorem˜3  This result immediately follows from Proposition˜14 by setting $\theta=0$, in which case